You are here: Foswiki>Commissioning Web>Logs>08Sep2012_Log (10 Sep 2012, JuanGuerra)Edit Attach

UT 20120908 AO#2 Commissioning

Observer: R. Bruglia, A. Puglisi, E. Pinna (LBTO)
Telescope Operator: D. Gonzalez(LBTO)
Additional AO Support: JCGuerra (LBTO)
Telescope Support: (LBTO)
Instrument Support: J.Morris (LBTO)

Summary

Installation of Retreo_Reflector. Alignment of the 4D. Alignment of the W#.

Acquisition and analysis of the modes using the 4D to compare with the modes recorded before the ASM electronic repair.

Set-up of the W# devices to work in the LFBG focal station and test closing the loop.

Hand-over 18:00

Details

4D.

Interferometer working properly, the alignment with the retro-reflector and secondary took a few hours. With the collimation beam the lase spot can be seeing close in the centre of M3 and looking to the Retro-reflector the spot was in the border of the back of the RR. Moving M3 we could centre the spot in the centre of the RR back face.

Positions: M3: -1135, 25, 0 -48.3

The beam expander is installed to illuminate the secondary. Checking the return image and twiking the hexapod the secondary image is full display in the interferometer image. The hexapod position is found at:

Hexapod position: 4.9, -1.22, 4.45, 590, 650 ,0

Secondary

Runa has acquired multiples interferograms images to analyse the status of the secondary actuators and if the model base acquired before the repair of the siggen is good. A report will be written when the data are analysed.

Looking the images of the interferogram and removing 50 modes( convection contribution and trifoild aberrations), the flat seems good with a RMS of 30nm.

A second set of data acquisitions will be tomorrow.

Left W# unit

Power on of the Unit, the software talk with the unit and to secondary through the fastlink. The system is aligned for the LFBG focal station moving the haxapod.

Hexapod position: 2.53,1.65,4.35,95,207,0is thinking to perform tomorrow.

A new rotation offset is determined and updated to the wfssx configuration files.

The IRTC performance ok and the image of the spot to analysise the psf is the right one. The spot is center in the center of the small field.

The AOS configuration file is updated with the LFBG focal station.

A close loop test is performed to check the system. It was done with a reconstructor with 400 modes and low gain. The AO system keep the loop close and we didn't perform any quality performances check of the system.

-- JuanGuerra - 09 Sep 2012
Topic revision: r4 - 10 Sep 2012, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback